OpenStudy (anonymous):

For (a) use a calculator to evaluate to four significant digits. Use the answer in (a) to evaluate (b) and (c).

OpenStudy (anonymous):

A\[\sin(7.3\pi)\]B\[\sin(20000000007.3\pi)\]C\[\cos(\frac{ \pi }{ 2 }-20000000007.3\pi)\]I got -.0890 for A, -.7400 for B, and .7400 for C. Am I correct?

OpenStudy (anonymous):

I get something else for b

OpenStudy (anonymous):

How did you get that for B and C?

OpenStudy (anonymous):

I did the first time too. I calculated it the first time as \[\sin(20000000007.3\pi)\]and punched it into my calculator and got -.8092, but then I got -.7400 when I calculated it like this \[\sin(2000000000)(-.8090)\]I don't know which is correct.

OpenStudy (anonymous):

The second one I actually use the answer I got in (a), but on the first one I didn't

OpenStudy (anonymous):

I don't see why you would multiply both A and B together.

OpenStudy (anonymous):

I don't know how else i'm suppose to use the answer of (a) in (b) and (c). Do you?

OpenStudy (anonymous):

Just compare answers I suppose.

OpenStudy (anonymous):

If I just punch it into my calculator, then I get -.8092 for (b) and -.8091 for (c)